OpenStudy (technodynamic):

Calculate the percent abundance. A fictitious element, deaborinium (Db), is made up of two isotopes. A sample of the element contains 941 atoms of ^292Db and 2923 atoms of ^295Db. Assuming that this is a statistically representative sample of the element.

OpenStudy (aaronq):

\(\sf percent ~abundance ~of~A=\dfrac{atoms~of~A}{Total~number~of~atoms}*100\%\) e.g. \(\sf percent ~abundance ~of~^{292}Db=\dfrac{atoms~of~^{292}Db}{^{292}Db~atoms+^{295}Db~atoms}*100\%\) \(\sf percent ~abundance ~of~^{292}Db=\dfrac{ 941}{ 941+2923}*100\%=24.353\%\)